OpenStudy (anonymous):

3x-y-z=3 -2x+3y+z=-3 4x-2y-2z=2 not understanding this

OpenStudy (anonymous):

are you familiar with the elimination method ?

OpenStudy (uri):

Cramer's rule?

OpenStudy (anonymous):

errrrm no

OpenStudy (anonymous):

i dont know what elimantion methid is

OpenStudy (uri):

Has they mentioned in which method you have to do this? @openstudy2222222

OpenStudy (anonymous):

think its elimantion methid but idk how to do that @Ⓐ uri

OpenStudy (anonymous):

take the first two equations and add them...can you do that ?

OpenStudy (anonymous):

the z's will cancel out

OpenStudy (anonymous):

yes ^_^ @kelliegirl33

OpenStudy (anonymous):

add them and let me see what you get

OpenStudy (anonymous):

i got 36 :\

OpenStudy (anonymous):

no.....let me show you....hold on

OpenStudy (anonymous):

3x - y - z = 3 -2x + 3y + z = -3 ----------------add x + 2y = 0 do you see how I did that ?

OpenStudy (anonymous):

OH i see :)

OpenStudy (anonymous):

now we have to take the last two equations and cancel out the z's. Sometimes when doing this we have to multiply an equation by a number to make the z's opposite...one positive and one negative with the same coefficient. I understand how this can be confusing so I will show you...hold on

OpenStudy (anonymous):

-2x + 3y + z = -3 -->(2)-2x + 3y + z = -3 4x - 2y - 2z = 2 ---------------- -4x + 6y + 2z = -6 (result of multiplying by 2) 4x - 2y -2z = 2 ---------------- wait.....there is a problem....the x's are cancelling out also...let me try something....give me a second

OpenStudy (anonymous):

Let me try to eliminate the y's 3x - y - z = 3 -->(3)3x - y - z = 3 -2x + 3y + z = -3 ---------------- 9x - 3y - 3z = 9 (result of multiplying by 3) -2x + 3y + z = -3 -----------------add 7x - 2z = 6 now take the last 2 equations and eliminate the y's -2x + 3y + z = -3 -->(2)-2x + 3y + z = -3 4x - 2y - 2z = 2 -->(3)4x - 2y - 2z = 2 -------------------- -4x + 6y + 2z = -6 (result of multiplying by 2) 12x - 6y -6z = 6 (result of multiplying by 3) --------------------add 8x - 4z = 0 I have eliminated the y's. I want you to look over this and ask any question about anything you do not understand. Take your time...I will wait

OpenStudy (anonymous):

okay so to elimate the y's do you have to multiply the top i still lost can yu explain it a lil more :[

OpenStudy (anonymous):

The elimination method...you have to eliminate a variable....either x,y,or z. I have just eliminated the y's. To do this there are times when you have to multiply to get the variables to cancel out. For instance, -4x and 4x will cancel out the x's. But if you have -2x and 3x, you would have to multiply one equation by 3 and one equation by 2 and then you would get -6x and 6x and they would cancel. Sometimes you have to multiply both equations, sometimes just 1 and sometimes none. Is this making any sense to you ?

OpenStudy (anonymous):

why did you log off......when you decide to get back on, tag me, and we will finish this problem.